# A Family of Berndt-Type Integrals and Associated Barnes Multiple Zeta Functions

### v1

@@ 4 Berndt-Type Integrals via Hyperbolic Series
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}+(-i)^{p+1}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{2^{p+1}}ip((1-i))^{p-1}\pi^{p}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p-1}}{\sinh\left((2n-1)\pi/2\right)\cosh^{2}\left((2n-1)\pi/2\right)}}$
$\quad+\displaystyle\frac{1}{8}((1-i))^{p+1}\pi^{p+1}\displaystyle\sum_{m=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{m}m^{p}}{\sinh(m\pi)\cosh^{2}(m\pi)}}-\displaystyle\frac{1}{2^{p+1}}i(1-i)^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\cosh^{3}\left((2n-1)\pi/2\right)}}$
$\quad-\displaystyle\frac{1}{2^{p+2}}i((1-i))^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\sinh^{2}\left((2n-1)\pi/2\right)\cosh\left((2n-1)\pi/2\right)}}.$ (4.21)
@@
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq argz\leq 0)$, and the segment $[-iR,0]$, the contour is illustrated below:
Obviously, there exist poles in case that
@@
This proof relies on analyzing the functions $Z_{i}(s,\theta;a,b)$($j=1,2,3,4$) and the usual residue computation. These functions are meromorphic in the entire complex plane with some simple poles at specific points.
First, we note that function $Z_{1}(s,\theta;a,b)$ has poles at $s=\pm\displaystyle\frac{n}{a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{n}{a}$ and by using Lemma 3.3, the residue is
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{n}{a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh(n\theta/a)}{\sinh(bn\pi/a)\cosh^{2}(bn\pi/a)}.$ (4.28)
As $s=\pm\displaystyle\frac{ni}{b}$ and the residue is
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin(n\theta/a)}{\sinh(an\pi/b)}.$ (4.29)
From Lemma 3.3, for all $n\in\mathbb{Z}$ we may deduce the asymptotic expansions of a few reciprocal quadratic trigonometric and hyperbolic functions as follows
@@
$\quad+o\left(\left(bs-\displaystyle\frac{2n-1}{2}i\right)^{2}\right)\quad\text{as }s\rightarrow\displaystyle\frac{2n-1}{2b}i.$ (4.30)
As $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ and the residue is
@@
$\quad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))\cosh((2n-1)a\pi/(2b))}{\sinh^{2}((2n-1)a\pi/(2b))}.$ (4.31)
Furthermore, applying (3.14) and (3.15) we deduce the expansion
$Z_{1}(s,\theta;a,b)=\displaystyle\frac{\pi^{2}}{\cosh^{2}(b\pi s)}\left(\displaystyle\frac{\theta}{ab}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6ab}s-\left(\displaystyle\frac{b}{a}-\displaystyle\frac{a}{b}\right)\zeta(2)\theta s+o(1)\right),\quad s\to 0.$
Therefore, as $s=0$ and the residue is
@@
By Lemma 3.1, summing the five contributions (4.28)-(4.32) yields the desired result (4.25).
Secondly, The function $Z_{2}(s,\theta;a,b)$ possesses poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{2n-1}{2a}$ and by applying Lemma 3.3, the residue is
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\cosh((2n-1)\theta/(2a))}{\sinh((2n-1)b\pi/(2a))\cosh^{2}((2n-1)b\pi/(2a))}.$ (4.33)
As $s=\pm\displaystyle\frac{ni}{b}$ and the residue is
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}.$ (4.34)
As $s=\pm\displaystyle\frac{(2n-1)i}{2b}$, for this double pole, we employ the expansion from (4) to obtain
@@
$\quad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))\sinh((2n-1)a\pi/(2b))}{\cosh^{2}((2n-1)a\pi/(2b))}.$ (4.35)
As $s=0$, using the expansions from (3.15) and (3.16), we derive
$Z_{2}(s,\theta;a,b)=\displaystyle\frac{\pi^{3}\cosh(\theta s)}{\cos(a\pi s)\cosh^{2}(b\pi s)}\left(\displaystyle\frac{1}{b}\displaystyle\frac{1}{s}-\zeta(2)bs+o(1)\right),\quad s\to 0,$
@@
At last, the function $Z_{3}(s,\theta;a,b)$ has poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$.
As $s=\pm\displaystyle\frac{2n-1}{2a}$ and by using Lemma 3.3, we compute the residue at these simple poles
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh((2n-1)\theta/(2a))}{\sinh^{2}((2n-1)b\pi/(2a))\cosh((2n-1)b\pi/(2a))}.$ (4.37)
As =±$=\pm\displaystyle\frac{(2n-1)i}{2b}$, the simple poles contribute
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{(2n-1)i}{2b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))}.$ (4.38)
As $s=\pm\displaystyle\frac{ni}{b}$, these double poles, we apply the asymptotic expansion from Lemma 3.3
$\left(\displaystyle\frac{\pi}{\sinh(b\pi s)}\right)^{2}$ $=\displaystyle\frac{1}{\left(bs-ni\right)^{2}}-2\zeta(2)+6\zeta(4)\left(bs-ni\right)^{2}+o\left(\left(bs-ni\right)^{2}\right)\quad\text{as }s\rightarrow ni,$
yielding
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{ni}{b}\right]$ $=\displaystyle\frac{\pi^{2}\theta}{b^{2}}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}$ (4.39)
$\quad-\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin(n\theta/b)\sinh(an\pi/b)}{\cosh^{2}(an\pi/b)}.$
As $s=0$, the asymptotic behavior near the origin is given by
$Z_{3}(s,\theta;a,b)=\displaystyle\frac{\pi^{2}}{\cos(a\pi s)\cosh(b\pi s)}\left(\displaystyle\frac{\theta}{b^{2}}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6b^{2}}s-2\zeta(2)\theta s+o(1)\right),\quad s\to 0,$
@@
Applying Lemma 3.1 to the sum of these four residue contributions (4.37)-(4.40) establishes identity (4.27).
@@
We prove the first identity in detail, as the proofs for the other two follow similarly. By taking the $p$-th odd derivative of (4.25) in Theorem 4.2 with respect to $\theta$, then $\theta\rightarrow 0$, we obtain the result as follows
$\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sinh\left(\displaystyle\frac{n\theta}{a}\right)\right]_{|\theta=0}=\left(\displaystyle\frac{n}{a}\right)^{p},\quad\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\theta\cos\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=p(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p-1},$
$\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sin\left(\displaystyle\frac{n\theta}{b}\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{n}{b}\right)^{p},\quad\displaystyle\frac{d^{p}}{d\theta^{p}}\left[\sin\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p}.$
@@
For $p=1$, we obtain the special case
@@
and for $p\geq 3$, we obtain the equation (4.41) .
The proofs of equations (4.42)and (4.43) follow similarly, with the following key differences: For (4.42), we take the $(p-1)$-th order derivative (i.e., even-order derivative) of equation (4.26) with respect to $\theta$; For (4.43), we take the $p$-th order odd derivative of equation (4.27) with respect to $\theta$. ∎
@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$G_{p,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{2}(ny)},$ (5.46)
$G_{p,2}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)},$ (5.47)
$\overline{G}_{p,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{2}((2n-1)y/2)\cosh((2n-1)y/2)}.$ (5.48)
@@
$G_{p,2}(y)=$ $(-1)^{(p-1)/2}\displaystyle\frac{(p-1)!z^{p+1}z^{\prime}x(1-x)}{2^{p+1}}\mathrm{R}_{p-1}(1-x)-\displaystyle\frac{pz^{p+1}z^{\prime}x(1-x)}{2^{p}}\mathrm{S}_{p-1}(1-x)\sqrt{1-x}$
@@
where $\mathrm{R}_{p-1}(1-x)=\displaystyle\frac{(-x)^{(p-3)/2}}{(p-1)!}q_{p-1}\left(\displaystyle\frac{1-x}{-x}\right)\in\mathbb{Q}[x]$ and ${q}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{sn}(u)^{2}$; ${\rm S}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{cd}\left(u\right)$.
@@
Beginning with equation (4.41), we derive
$G_{p,2}(y)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}X_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{p\pi^{p}}{2^{p-1}y^{p+1}}Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+2}}{2^{p}y^{p+2}}\text{DY}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.50)
Now, we need to compute the series $Y_{p,1}\left(y\right)$. According to Lemma 3.5, we obtain the Maclaurin expansion for ${\rm cd}(u)$
@@
And in the page $165$ of Ramanujan’s Notebooks (III)[5], we get
@@
Applying $q\equiv q\left(x\right)=e^{-y}$, then
${\rm cd}(u)$ $=\displaystyle\frac{2\pi}{\mathrm{K}k}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{\mathrm{K}k}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}Y_{2j+1,1}\left(y\right).$ (5.54)

### v3

@@ 4 Berndt-Type Integrals via Hyperbolic Series
$\left(1+(-i)^{p+1}\right)\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{p}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{2^{p+1}}ip(1-i)^{p-1}\pi^{p}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p-1}}{\sinh\left((2n-1)\pi/2\right)\cosh^{2}\left((2n-1)\pi/2\right)}}$
$\quad+\displaystyle\frac{1}{8}(1-i)^{p+1}\pi^{p+1}\displaystyle\sum_{m=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{m}m^{p}}{\sinh(m\pi)\cosh^{2}(m\pi)}}-\displaystyle\frac{1}{2^{p+1}}i(1-i)^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\cosh^{3}\left((2n-1)\pi/2\right)}}$
$\quad-\displaystyle\frac{1}{2^{p+2}}i(1-i)^{p-1}\pi^{p+1}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n}(2n-1)^{p}}{\sinh^{2}\left((2n-1)\pi/2\right)\cosh\left((2n-1)\pi/2\right)}}.$ (4.21)
@@
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the quarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq argz\leq 0)$, and the segment [iR,0][-iR,0], the contour is illustrated below:
The poles are located at
@@
This proof relies on analyzing some functions and the usual residue computation. These functions are meromorphic in the entire complex plane with some simple poles at specific points.
First, we note that function $Z_{1}:=Z(s,\theta;a,b|1,0,1,2)$ has poles at $s=\pm\displaystyle\frac{n}{a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles) and $s=0$ (simple poles), where $n\in\mathbb{N}$.
By using Lemma 3.3, for all $n\in\mathbb{Z}$ we may deduce the asymptotic expansions of a few reciprocal quadratic trigonometric and hyperbolic functions as follows
@@
$\quad+o\left(\left(bs-\displaystyle\frac{2n-1}{2}i\right)^{2}\right).$ (4.28)
After direct computation, the residues of $Z_{1}$ at the simple poles $\pm\displaystyle\frac{n}{a}$ and $s=\pm\displaystyle\frac{ni}{b}$, as well as the residues at the second-order poles $\pm\displaystyle\frac{(2n-1)i}{2b}$ derived by applying (4), are respectively as follows:
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{n}{a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh(n\theta/a)}{\sinh(bn\pi/a)\cosh^{2}(bn\pi/a)},$ (4.29)
$\mathrm{Res}\left[Z_{1},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin(n\theta/a)}{\sinh(an\pi/b)},$ (4.30)
@@
$\qquad\qquad\qquad\qquad\qquad\qquad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))\cosh((2n-1)a\pi/(2b))}{\sinh^{2}((2n-1)a\pi/(2b))}.$ (4.31)
Furthermore, by applying equations (3.14) and (3.15), along with the power series expansion of the hyperbolic sine function, we deduce the following expansion:
$Z_{1}=\displaystyle\frac{\pi^{2}}{\cosh^{2}(b\pi s)}\left(\displaystyle\frac{\theta}{ab}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6ab}s-\left(\displaystyle\frac{b}{a}-\displaystyle\frac{a}{b}\right)\zeta(2)\theta s+o(1)\right),\quad s\to 0.$
Then, using the above expansion, we can compute the residue of $Z_{1}$ at 0 as
@@
By Lemma 3.1, summing the four contributions (4.29)-(4.32) yields the desired result (4.25).
Secondly, the function $Z_{2}:=\bar{Z}(s,\theta;a,b|0,1,1,2)$ possesses poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (double poles) and $s=0$ (simple poles), where $n\in\mathbb{N}$. Through direct computation, we obtain the following residue values of $Z_{2}$ at the simple poles $\pm\displaystyle\frac{2n-1}{2a}$ and $\pm\displaystyle\frac{ni}{b}$, as well as at the second-order poles $\pm\displaystyle\frac{(2n-1)i}{2b}$:
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\cosh((2n-1)\theta/(2a))}{\sinh((2n-1)b\pi/(2a))\cosh^{2}((2n-1)b\pi/(2a))},$ (4.33)
$\mathrm{Res}\left[Z_{2},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)},$ (4.34)
@@
$\qquad\qquad\qquad\qquad\qquad+\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\cos((2n-1)\theta/(2b))\sinh((2n-1)a\pi/(2b))}{\cosh^{2}((2n-1)a\pi/(2b))}.$ (4.35)
When $s=0$, using the expansions from (3.15) and (3.16), we derive
$Z_{2}(s,\theta;a,b)=\displaystyle\frac{\pi^{3}\cosh(\theta s)}{\cos(a\pi s)\cosh^{2}(b\pi s)}\left(\displaystyle\frac{1}{bs}-\zeta(2)bs+o(1)\right),\quad s\to 0,$
@@
At last, the function $Z_{3}:=Z(s,\theta;a,b|0,1,2,1)$ admits poles at $s=\pm\displaystyle\frac{2n-1}{2a}$ (simple poles), $s=\pm\displaystyle\frac{(2n-1)i}{2b}$ (simple poles), $s=\pm\displaystyle\frac{ni}{b}$ (double poles), $s=0$ (simple poles), where $n\in\mathbb{N}$. We apply the expansion given in Lemma 3.3 to obtain
$\left(\displaystyle\frac{\pi}{\sinh(b\pi s)}\right)^{2}$ $=\displaystyle\frac{1}{\left(bs-ni\right)^{2}}-2\zeta(2)+6\zeta(4)\left(bs-ni\right)^{2}+o\left(\left(bs-ni\right)^{2}\right)\quad\text{as }s\rightarrow ni.$ (4.37)
Therefore, by performing direct calculations and applying (4.37), the residues of $Z_{3}$ at the simple poles $\pm\displaystyle\frac{2n-1}{2a}$ and $\pm\displaystyle\frac{(2n-1)i}{2b}$, as well as at the second-order poles $\pm\displaystyle\frac{ni}{b}$, can be readily obtained as follows:
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{2n-1}{2a}\right]=\displaystyle\frac{\pi^{3}}{a}\displaystyle\frac{(-1)^{n}\sinh((2n-1)\theta/(2a))}{\sinh^{2}((2n-1)b\pi/(2a))\cosh((2n-1)b\pi/(2a))},$ (4.38)
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{(2n-1)i}{2b}\right]=\displaystyle\frac{\pi^{3}}{b}\displaystyle\frac{(-1)^{n}\sin((2n-1)\theta/(2b))}{\cosh((2n-1)a\pi/(2b))},$ (4.39)
$\mathrm{Res}\left[Z_{3},s=\pm\displaystyle\frac{ni}{b}\right]=\displaystyle\frac{\pi^{2}\theta}{b^{2}}\displaystyle\frac{(-1)^{n}\cos(n\theta/b)}{\cosh(an\pi/b)}-\displaystyle\frac{a\pi^{3}}{b^{2}}\displaystyle\frac{(-1)^{n}\sin(n\theta/b)\sinh(an\pi/b)}{\cosh^{2}(an\pi/b)}.$ (4.40)
Applying (4.37) (with $n=0$) and the power series expansion of the hyperbolic sine function, the following expansion of $Z_{3}$ at zero can be computed:
$Z_{3}=\displaystyle\frac{\pi^{2}}{\cos(a\pi s)\cosh(b\pi s)}\left(\displaystyle\frac{\theta}{b^{2}}\displaystyle\frac{1}{s}+\displaystyle\frac{\theta^{3}}{6b^{2}}s-2\zeta(2)\theta s+o(1)\right),\quad s\to 0,$
@@
Applying Lemma 3.1 to the total of four residue terms (4.38)-(4.41) establishes identity (4.27).
@@
We prove the first identity in detail, as the proofs for the other two follow similarly. By taking the $p$-th odd derivative of (4.25) in Theorem 4.2 with respect to $\theta$, then $\theta\rightarrow 0$, we acquire the result as follows
$\displaystyle\frac{\mathrm{d}^{p}}{\mathrm{d}\theta^{p}}\left[\sinh\left(\displaystyle\frac{n\theta}{a}\right)\right]_{|\theta=0}=\left(\displaystyle\frac{n}{a}\right)^{p},\quad\displaystyle\frac{\mathrm{d}^{p}}{\mathrm{d}\theta^{p}}\left[\theta\cos\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=p(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p-1},$
$\displaystyle\frac{\mathrm{d}^{p}}{\mathrm{d}\theta^{p}}\left[\sin\left(\displaystyle\frac{n\theta}{b}\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{n}{b}\right)^{p},\quad\displaystyle\frac{\mathrm{d}^{p}}{\mathrm{d}\theta^{p}}\left[\sin\left(\displaystyle\frac{2n-1}{2b}\theta\right)\right]_{|\theta=0}=(-1)^{(p-1)/2}\left(\displaystyle\frac{2n-1}{2b}\right)^{p}.$
@@
For $p=1$, we derive the special case
@@
and for $p\geq 3$, we deduce (4.42) .
The proofs of (4.43) and (4.44) follow similarly, with the following key differences: For (4.43), we take the (p1)(p-1)-th order derivative (i.e., even-order derivative) of equation (4.26) with respect to $\theta$. For (4.44), we employ the $p$-th order odd derivative of equation (4.27) with respect to $\theta$. ∎
@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$G_{p,1,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{2}(ny)},$ (5.47)
$\tilde{G}_{p,1,2}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh((2n-1)y/2)\cosh^{2}((2n-1)y/2)},$ (5.48)
$\tilde{G}_{p,2,1}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{2}((2n-1)y/2)\cosh((2n-1)y/2)}$ (5.49)
@@
$G_{p,1,2}(y)=$ $(-1)^{(p-1)/2}\displaystyle\frac{(p-1)!z^{p+1}z^{\prime}x(1-x)}{2^{p+1}}\mathrm{R}_{p-1}(1-x)-\displaystyle\frac{pz^{p+1}z^{\prime}x(1-x)}{2^{p}}\mathrm{S}_{p-1}(1-x)\sqrt{1-x}$
@@
where $\mathrm{R}_{p-1}(1-x)=\displaystyle\frac{(-x)^{(p-3)/2}}{(p-1)!}Q_{p-1}\left(\displaystyle\frac{1-x}{-x}\right)\in\mathbb{Q}[x]$ and ${Q}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{sn}^{2}(u)$; ${\rm S}_{p-1}\left(x\right)$ represents coefficients in the Maclaurin expansion of $\mathrm{cd}\left(u\right)$.
@@
Beginning with (4.42), we derive
$G_{p,1,2}(y)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}G_{p,1,0}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{p\pi^{p}}{2^{p-1}y^{p+1}}\tilde{G}_{p-1,1,0}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+2}}{2^{p}y^{p+2}}\tilde{G}_{p,2,-1}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.51)
Now, we need to compute the series $\tilde{G}_{p-1,1,0}\left(y\right)$. According to Lemma 3.5, we obtain the Maclaurin expansion for ${\rm cd}(u)$
@@
Applying $q\equiv q\left(x\right):=e^{-y}$ to (5.53), we obtain
${\rm cd}(u)$ $=\displaystyle\frac{2\pi}{{K}k}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{{K}k}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2{K}}\right)^{2j}}{\left(2j\right)!}}\tilde{G}_{2j,1,0}\left(y\right).$ (5.54)
And in the p. $165$ of Ramanujan’s Notebooks (III)[5], we obtain